Please help me answer this supply chain question, i will rate. thanks!!

Considor the perfect order caloulation for Bartey Company Last year Bartley Company experiencedthe followng resutb artley Company. Last year Bartley 6.2 million orders processed the following 25,000 orders delivered late 15,000 orders incomplete 10,000 orders damaged 10,000 orders billed incorrectly Furthermore, these 60,00 ailures were spread across 30,000 orders, which meant that some orders had more than one problem. Calculate the percentage of perfect orders # all performance results above remained the same excespt a. 10,000 are delivered late, and total failures are now spread across 45,000 orders. The percentage of perfect orders is ?% if 10,000 are delivered late, and total failures are spread across 45,000 orders. (Enter b. 10,000 are delivered late, but total failures are still spread across 30,000 orders. y esponse rounded to e do rnal place The percentage of perfect orders is10,000 are delivered late, but total failures are stll pread across 30,000 orders.(Enter your response rounded to one decimal place) c. According to the logic of the perfect order measure, does an incorrectly billed order have the same impact as a damaged order? O A. An incorrectly billed order has a lower impact than a damaged order O B. An incorrectly billed order has the same impact as a damaged order O C. An incorrectly billed order has a higher impact than a damaged order O D. It is impossible to give the answer without additional information.
Answer #1

The perfect order rate calculation has a compounding formula where all these metrics are measured as the percentage of the preceding metric. But since the total number of failures is already mentioned in both the parts, we do not need to use it. Also since every sort of failure equally impacts the accuracy, so it does not matter how much is the late delivery rate. It will still be counted as a failure if it has only one defect.

a). % of perfect orders = (total orders - defective orders)*100/Total orders = (6200000 - 45000)*100/6200000 = 99.3%

b). % of perfect orders = (total orders - defective orders)*100/Total orders = (6200000 - 30000)*100/6200000 = 99.5%

c). Correct answer is B as both the defects have the same impact